Circular No. 7232

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1999 P1 (MACHHOLZ 2)

Ephemeris from the orbital elements on IAUC 7231:

1999 TT α_{2000} δ_{2000} Δ r ϵ β m_{2}
July 31 16^{h}52.^{m}54 −11^{o}22.^{′}0 1.226 1.990 125.^{o}0 24.^{o}7 20.6
Aug. 10 16 44.79 −10 56.1 1.237 1.888 113.7 29.4 20.6
20 16 41.28 −10 43.6 1.253 1.784 103.4 33.5 20.6
30 16 41.99 −10 42.8 1.268 1.677  94.2 36.9 20.6
Sept. 9 16 46.75 −10 50.9 1.277 1.568  85.8 39.8 20.6
19 16 55.37 −11 04.9 1.276 1.457  78.4 42.5 20.5
29 17 07.65 −11 21.5 1.261 1.345  71.9 45.1 20.4
Oct. 9 17 23.50 −11 37.2 1.231 1.232  66.1 47.9 20.2
19 17 42.94 −11 49.0 1.182 1.121  61.2 51.2 20.0
29 18 06.01 −11 54.0 1.114 1.013  57.1 55.4 19.8

Z. Sekanina (1999, A & Ap 342, 285; Table 7) tabulates the expected offsets of components B and D from component A. In terms of ΔT, these amount to +0.21 and +0.82 day, respectively.

COMET C/1999 H1 (LEE) Total visual magnitudes: June 22.27 UT, 6.6 (S. T. Rae, Palmerston

North, New Zealand, 10 × 50 binoculars); July 3.26, 6.8 (N. Biver, Oahu, HI, 0.26-m reflector) 27.09, 6.1: (J. Shanklin, Cambridge, England, 14 × 100 binoculars); 29.09, 6.6 (K. Hornoch, Lelekovice, Czech Republic, 10 × 50 binoculars); Aug. 1.01, 6.6 (B. H. Granslo, Fjellhamar, Norway, 0.07m refractor); 2.07, 6.8 (R. Bouma, Groningen, The Netherlands, 0.16-m reflector); 2.35, 7.2 (J. Bortle, Stormville, NY, 20 × 80 binoculars).

V382 VELORUM A. C. Gilmore, Mount John University Observatory, provides further

photometric observations, continuing the series on IAUC 7226: July 20.476 UT, V = 7.73, U − B = −0.93, B − V = −0.23, V − R = +1.23, V − I = +0.55, airmass = 2.92; 22.329, 7.71, −0.92, −0.24, +1.17, +0.52, 1.39. Standard deviations < 0.02 mag on July 20 and < 0.01 mag on July 22.

V1493 AQUILAE Visual magnitudes: July 17.92 UT, 11.4 (R. Bouma, Lattrop, The

Netherlands); 21.96, 12.4 (M. Reszelski, Szamotuly, Poland); 26.08, 12.3 (P. Schmeer, Bischmisheim, Germany); 29.85, 12.8 K. Hornoch, Lelekovice, Czech Rep.); Aug. 3.89, 13.0 (M. Lehký, Hradec Kralové, Czech Rep.).

1999 August 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Brian G. Marsden